Multiply and simplify:
(-5√6)(2√3)

Answer by nyc_function(2741)

(-5√6)(2√3) = -10(√18)
-10(√18) = 10(√9(√2))
10(√9(√2)) = 10(3)(√2) = 30√2
Done!